Use words or set notation to identify equations that have no solution, or equations that are true for all real numbers.
5x-5=3x-7+2(x+1)

Answer by stanbon(75887)

5x-5=3x-7+2(x+1)
5x-5 = 3x -7 + 2x + 2
5x-5 = 5x-5
---
This is true for all values of x.
=====================================
Cheers,
Stan H.